OpenStudy (anonymous):

Simplify the expression..fourth root of 8 muliplied by the fourth root of 8

OpenStudy (pfenn1):

\[\sqrt[4]{8}\sqrt[4]{8}=8^{\frac14}8^{\frac14}\]Remember \[x^ax^b=x^{a+b}\]